OpenStudy (anonymous):

solve 2x^2+(x+12)(x-4)=3(x-8)^2

OpenStudy (anonymous):

multiply out, get a quadratic equation

OpenStudy (anonymous):

actually i lied. multiply out, get a linear equation!

OpenStudy (anonymous):

I didn't even get a quadratic, I got a linear.

OpenStudy (anonymous):

\[2x^{2}+x^{2}+8x-48=3(x^{2}-16x+64)\]\[3x^{2}+8x-48=3x^{2}-48x+192\]\[56x=240\]\[7x=30\]\[x=30/7\]